OpenStudy (anonymous):

Use Euler’s Formula to find the missing number. Vertices: 10 Edges: 29 Faces: ? (1 point) 22 21 24 20

OpenStudy (campbell_st):

the formula is V + F = E + 2 you have V = 10 and E = 29 then 10 + F = 29 + 2 or 10 + F = 31 fnow solve for F

OpenStudy (anonymous):

21